Grade 12Algebra

The vapour pressure of pure water at 30°C is 31.80 mm of hg. How many grams of urea(molecular mass =60) should be dissolved in 1000 g of water to lower the vapour pressure by 0.25 mm of hg?

khushbu
9 Years agoGrade 12

3 Answers

Manan patidar
8 Years ago
The vapour pressure of pure water at 30°C is 31.80 mm of hg. How many grams of urea(molecular mass =60) should be dissolved in 1000 g of water to lower the vapour pressure by 0.25 mm of hg?
Answer is 26.2g
 
Arun
8 Years ago
Dear student
 
 
As we know
 
p° - p /p° = W2 M1 /M2 W1
 
Now put the values
 
0.25 /31.8 = W2 *18 /(60* 100)
 
On calculation,
 
W2 = 2.62 gm
 
 
Regards
Arun (askIITians forum expert)
Laishngbam karoke
5 Years ago
Given
 Urea (A), WA= ?
                 MA= 60g/mol
 
Water (B), WB = 100g
                   MB = 18g/mol
P° = 31.80mm
P°- Ps = 0.25mm
Now, 
         P° - Ps /P° = WA /MA× MB / WB
         0.25/31.80= WA /60× 18 /100
          0.25/31.80= WA/30 × 9/100
         
          WA = 0.25×30× 100/31.80× 9
                  = 0.25× 10 × 100/31.80× 3
                   = 250/31.8 ×3 
                   = 2500/954
     WA        = 26.2 g